NAOJ GW Elog Logbook 3.2

By using TAMA demodulator, I monitor its output with two identical RF frequency signals as inputs. The signal drifts from 83.1 to 81.9.

Marc, Michael, and Yuhang
When we lock filter cavity with GR, IR detuning has change related to alignment. When GR automatic alignment (AA) and pointing loop is closed, IR detuning change can be stabilized.The filter cavity pointing loop is working mainly to fix the injection beam on end mirror. AA works to align filter cavity to the incident beam.
To check how detuning change for different alignment condition, we can change the pointing. By pointing the incident beam to different positions on end mirror and keeping AA loop closed, we can get AOM frequency for each point on end mirror when BAB is on resonance. The change of position on end mirror gives us a dependence of detuning as end mirror beam hitting point. In this way, we call it a detuning map for end mirror.
The changed parameters are not only beam hitting position on end mirror, the other changed parameters are input/end mirror angles and cavity length. A typical change for input angle is 40urad, end angle is 10urad and cavity length is 0.2um. Since the optical axis of filter cavity is almost the same for GR and IR, the GR AA should work also for IR. In addition, GR length control should also work for IR. Therefore, the map we get should just depend on beam hitting position of end mirror. The corresponding map is attached.

Today, I went to filter cavity end room and centered the FC GR transmission camera in yaw. But I didn't move pitch.
The pointing offset is still a good reference, since I didn't move PSD.

The filter cavity IR detuning is monitored from 2021-07-08 11:40 to 2021-07-09 11:40. (The time used in DGS is JST minus 9 hours) The minute trend data is saved in standalone desktop/detuning/20210709.
The screenshot of this monitor is attached.
Although there are many peaks in the detuning data, only 4 of them come from the unlock of filter cavity. Others are due to the suspended mirror sudden position changes but pointing loop has limited bandwidth.
We see change of detuning even when fc length is controlled.
The FC length control change may also come from the main laser frequency change, which is due to we use laser frequency as a reference at low frequency. Especially, we don't have a reference cavity as used in gravitational wave detectors, such as KAGRA.

Abe-san, Aso-san, Marc, Michael
For reference, ETMY cleaning is summarized in KLOG entries : 17219, 17271, 17292, 17311, 17397, 17409
We brought the ETMY inside PCI room using the crane.
We added HEPA filters and put an ion gun at the end of the pressured air.
We remove optics from the small optical table on the side of the PCI setup and installed ETMY box on it.
Using a strong green light and a strong white flashlight we inspected and slightly cleaned the AR surface using ultra pure water and the ion gun.
On this table, we installed ETMY inside its holder using 4 jacks.
In order to avoid incident, we removed the entire imaging unit optical table to ease the ETMY installation on the translation stage.
Before doing so, we had installed pairs of forks on 3 of the 4 pillars to be sure to recover the same imaging unit position.
We installed ETMY on the translation stage (the additional weight due to the jig is negligeable because the translation stage can hold few hundreds kg).
We removed the HR surface first contact while using the ion gun.
To avoid scratching ETMY surface or magnets, we decided to let a metal ring at the edge of the mirror surface.
We reinstalled the IU and turning on the probe beam showed that this beam was still hitting well on the IU optics.
I don't have so much pictures but we'll add beautiful ones to this entry.
This morning we set up translation stage limit along the Z axis. We are now not letting ETMY get closer than 1 or 2 cm to both side.
There are now really strange troubles with Zaber that does not recognize the translation stage (or any com port) even if it is working fine in labview..
We will solve this issue before any translation stage motion.

Here is the calibration performed before the installation of ETMY on PCI setup :
AC_surfref = 0.45825;
DC_surfref = 3.987;
P_in = 30.3e-3;
abs_surfref = 0.22;
R_surf = 17.24 /W
The AC peak is located at Z=39.6 mm.

It was noticed that filter cavity z-correction was feeding back some high frequency components recently.
I modified the filters and now less high frequency components are feed to end mirror.
The new filter is called dc_damp2 (gain is adjusted so that we can use gain 1 in medm). Let's use this filter in the future.
Poles: 1e-4, 0.1(1), 20, 20, 300
Zeros: 0.04, 0.05(1), 3
The comparison of signal sent to end mirror is attached.

Marc and Yuhang
Recently, we found a new spot of filter cavity elog2573, which makes the IR locking accuracy much better (the spectrum below~3Hz reduce by up to a factor of 10). At the beginning, we thought we found a more stable optical axis. However, we did a test of GR length correction signal when using old and new spot, which shows pretty similar spectrum at frequency region below ~3Hz (attached figure 1). Since the GR length correction signal below ~3Hz tells us mirror motion information, this means the mirror motion is similar for the old and new spot.
Meanwhile, the correction measured in this time is different from elog2312. Especially, it seems more high frequency signal is sent to end mirror.

When we leave z-correction loop open, if we send 1000 excitation at 0.1Hz to channel "END-len-ex2", we get 32.4 counts from PZT mon(1620counts from figure REF4, which is amplified by 50), which means 3240 counts are sent to main laser (PZT mon is 100 times smaller than the signal sent to main laser PZT). 3240/1000*0.31=1V. It corresponds to 2MHz of main laser frequency change. It corresponds to 2um change in cavity length.
When we leave z-correction loop closed, if we send 1000 excitation at 0.1Hz to channel "END-len-ex2", we get 1934 counts from z-correction loop correction signal. (there are still 10 counts sent to PZT, but since it is so small, we neglect it). Therefore, 1934 counts corresponds to 2 um length correction.
For correction signal sent to end mirror, the calibration factor is thus 1.034 um/kcount.
(the data of this calibration is saved in standalone desktop/detuning/20210628)
In the last calibration calculation, I didn't consider the loop gain. Therefore, the calibration factor must have some error.
Nevertheless, we can use another way to do this calibration without considering the loop gain.
0. Lock filter cavity.
1. Change slightly the temperature of main laser.
2. Read how much main laser frequency is changed.
3. Check how much length correction is sent to end mirror.
I did these procedures. The frequency change is read from the attached two figures. The correction signal change is in the attached figure three.
And get calibration factor (frequency difference)/(correction signal) = (248.6-235.2) [MHz]/ (5200) [counts] = 2.56 [MHz] / 1000 [counts]
Since 1pm = 1Hz, we can calibrate the factor above as 2.56 [um]/[kcounts].

Because Pin ~3W seems quite convenient to measure viewport absorption (as it makes the 2 surfaces visible on the AC signal), we performed again the absorption measurement of the spare viewport.
Figure 1 shows a long Z scan of the translation stage with the 2 surfaces visibles and at the expected position.
Figure 2 and 3 show the absorption of respectively the spare and cleaned viewport.
The spare viewport shows some dusty spots (some visible by eye+ strong green light).
The cleaned viewport shows more dirty spots (also some visible by eye+ strong green light).
However, it seems that there is no more large stain pattern visible.

In the last week, resonant condition between GR and IR changed by around 20Hz.
To check if there is any correlation with suspended mirrors, I checked the oplev signal of all mirrors. Basically all mirrors are staying in the same orientiation, but end mirrors have quite obvious drift during the last week. In fact, this drift seems to be not really because we are not really correcting it by the coils. So we need to investigate why end mirror oplev is behaving like this.

The high frequency noise is same for old and new beam spots, but is increased for 50Hz detuning compared with the one on resonance. This noise difference could be explained by the cavity pole effect. The cavity pole effect for 50Hz detuning (half detune) is smaller than the one on resonance by a factor of ~sqrt(2). Please check P.50 of LIGO-T1800447 for the cavity pole effect of detuned cavity.

I measured FDS with CCFC with old beam spot. However, the result is similar to the one with new beam spot...
So the bump around 50Hz and larger detuning fluctuation will not be related to the beam spot position.
We had better FDS spectra before with old beam spot. I don't know why it is worse now...

I was using AOM scanning speed as 4000Hz/1.7s in the calibration. However, since the scanning speed for IR is 1/2 of the value for GR, the figure in the old elog was wrong.
Calibration for the measured spectrum should be: calibration = 2000/1.66666*11.5/11.2 #Hz/V (PDH: 11.2mV/11.5ms) (AOM: 4000Hz/1.66666s)
There was also problem for the calibration for off-center on-resonance, I modified the plot by using a more reasonable calibration. It comes from the center on-resonance. The new plot is shown in the attached figure.
We can see the new stable optical axis makes especially the low frequency length noise reduced. However, the high frequency noise is increased a bit.

Michael and Yuhang
To check the stability of detuning with stable cavity axis, GR AA and GR pointing, we decide to perform FDS measurement. To better resolve the whole rotation part of squeezing, we detuned filter cavity by 200Hz to avoid back scattered noise contamination.
The measurement result is in the attached figure. We can see the detuning fluctuation is about 10Hz from this measurement.
We still miss the homodyne angle around 90deg, we should take it soon.

Today I decided to further increase the laser power. Indeed my concern was coming from the fact that we could not clearly see the effects of the viewport 2 surfaces on the ac nor the phase signal.
I choose HWP angle = 55 degrees which translate to Pin = 3.135 W.
Figure 1 shows the results of a large Z scan of the translation stage : The 2 surfaces are now visibles (spikes in the AC/DC and phase jump) !
The surface with the smaller Z is the surface we want to measure (it is close to the expected value of 41.5 +30 mm ).
I checked the tilt of the viewport and it is still around 0.3 mm over the entire map area (30 mm radius).
A result of absorption measurement is presented in figure 2.
The strange thing is that it is really coherent with previous measurements....
On this measurement, one spot was saturating the AC signal so I started a new measurement with sensitivity 1 V (max) instead of 50 mV.
I really suspect 2 high absorption spots to be due to dust as it is quite visible by eye and seems different than the other drop like stains.
If this assumtion is correct (maybe can be checked with another measurement after applying first contact), it means that after cleaning, most of the remaining dirty things are mainly generating absorption below 100 ppm.

[Aritomi, Michael, Yuhang]
Today CC2 mass feedback was very unstable with gain of 2.7. We found that coil output to input mass was too large, so Yuhang offloaded the input mass with picomotor. After that, CC2 mass feedback becomes stable with gain of 2.7.
We tried to figure out the cause of 100Hz bump and found that the glitch appeared in CCFC error signal when we touched the SMA cable for CC1 mixer. We tightened it.
Then we measured FDS with CCFC (Fig 1). The degradation parameters are same as elog2597. The 100Hz bump maybe a bit better, but still present... The detuning fluctuation is ~20Hz. I noticed that the detuning seems anti-correlated to the homodyne angle.
To determine the squeezing level and generated squeezing precisely, I measured the shot noise and nonlinear gain just after the FDS measurement. The p pol PLL frequency for 20mW was 185MHz and BAB maximum was 282mV with 16mW power meter range. The BAB maximum without green is 56.8mV with 240MHz of p pol PLL frequency. This means the nonlinear gain is 5, which corresponds to the generated squeezing of 10.8dB.
For precise degradation budget, it is very important to measure the shot noise and nonlinear gain just after (before) the FDS measurement.

This measurement was done on 20210622.
I measured the CCFC locking accuracy with the new beam spot (Fig 1). CCFC calibration amplitude is 166mVpp and CCFC filter is gain of 1000 and LPF of 30Hz. CC2 mass feedback gain is 2.7.
Compared with old beam spot, the locking accuracy without CCFC is smaller. The locking accuracy with CCFC is similar to before because it is limited by high frequency noise.
Then I measured FDS with CCFC (Fig 2). The squeezing degradation parameters are as follows. Here are some points of this result.
- Because of improved reflection mode matching, the propagation loss is lower than before and squeezing level is 2.4dB. The propagation loss is 49%, which is consistent with 36% of propagation loss in PRL paper plus 20% pick off (0.64*0.8=0.51).
- In addition to the 100Hz bump, there is a large bump around 50Hz. This bump should be related to the new beam spot because this bump was not present in the old beam spot.
- The detuning fluctuation is ~20Hz with new beam spot. I think it is better to use the old beam spot for CCFC FDS measurement due to the bump around 50Hz and larger detuning fluctuation.
sqz_dB = 10.8; % generated squeezing
L_rt = 120e-6; % FC losses
L_inj = 0.32; % Injection losses
L_ro = 0.25; % Readout losses (propagation loss is 49%)
A0 = 0.06; % Squeezer/filter cavity mode mismatch
C0 = 0.02; % Squeezer/local oscillator mode mismatch
ERR_L = 1e-12; % Lock accuracy (m)
ERR_csi = 30e-3; % Phase noise (rad)

Abe, Marc
Following the Z scan of the cleaned viewport (see entry 2595) we decided to cross-check the relative alignment of the pump and probe laser using the surface reference sample.
We installed it and found out that it was quite misaligned (see figure 1)....
It means that over few days the alignment deteriorated quite a lot. We would really investigate how to get more stable alignment (maybe using similar mount provider as for the FC critical optics reported in entries 2583 2593)
We then spent most of the days trying to recover the proper alignment condition and finally reached the condition in figure 2.
To go from figure 1 to figure 2 we :
- checked the IU position (still 68 mm)
- checked the translation stage z position. By trying to have same amplitude of the lateral peaks in the AC signal we got z = 41.5 mm.
This gives us a new calibration factor :
ac = 0.4515 V
dc = 4.178 V
acdc = 0.108
p = 29.8 mW
R = 16.47 /W
Finally we reinstalled the cleaned viewport on the translation stage.

Abe, Marc
On Monday we decided to further increase the laser power by tuning the HWP.
In order to avoid burning dust and/or stains, we decided to do smaller map (radius = 9 mm) centered at X = 324 mm and Y = 134 mm.
Indeed, in this area no spikes were present in the previous measurement (see entry 2589).
The resulting map with Pin = 0.6236 W is presented in figure 1. In this measurement the lockin amplifier sensitivity was set to 500 uV.
There was some spikes saturating but we thought it could be good enough to estimate the viewport surface background absorption.
The resulting absorption is presented in figure 2. Again, the background level seems lower than the spare viewport one (see entry 2585). Another difference is the presence of many spikes (~2 times larger absorption than overall area).
This lower absorption could be explained by several possibilities (too low power to distinguish absorption, misalignment of probe and/or pump lasers, viewport tilted, properties of spare and cleaned viewport are differents, cleaning damaged the surface, ....)
To eliminate possibilities, we increased the power to Pin = 1.2503 W (HWP angle = 55 deg). and got the results reported in figure 3. The background absorption level stayed the same, meaning that we are indeed sensing absorption.
(note that the spike absorption level is not meaningful because the locking amplifier was saturating on purpose).
To eliminate the possibility of a viewport tilt, we then started to do Z scan to check the surface position across the viewport surface.
We changed back the laser power to Pin ~ 0.6W and did a Z scan at the top position of entry 2585. Result is presented in figure 4.
While the phase exhibits the expected behavior, I'm a bit more surprised by the AC and DC shapes...